# MS-0001-1922.65

### v1

@@ 1 Introduction
The medical device approval pathway is a function of the level of control necessary to provide reasonable assurance of a device’s safety and effectiveness (FDA 2018a). In general, devices posing a greater degree of risk are denoted by class and face greater regulatory controls. Although there are four common types of approval pathways (FDA 2022a), the vast majority of medical devices are cleared under the Premarket Notification 510(k) pathway. For the year 2022, over three thousand devices were cleared under the 510(k) pathway (FDA 2023a), while less than two hundred devices were cleared under the Humanitarian Device Exemption, Premarket Approval, and De Novo pathways combined (FDA 2023e, FDA 2023d, and FDA 2023c).
@@
The 510(k) pathway was developed both to reduce the burden for device manufacturers bringing medium-to-low risk (Class II and I) devices to market and to address the limited resources of the United States Food and Drug Administration (FDA) (Kramer and Yeh 2023). Devices cleared under the 510(k) pathway must demonstrate that the device is “substantially equivalent” to a legally marketed device, commonly referred to as a predicate device (FDA 2022b). A predicate device can be (a) any device that is legally marketed prior to May 28, 1976 (a preamendments device), for which clinical testing was not required, (b) any device that was reclassified from Class III (high risk) to Class II or Class I (medium-to-low-risk), or (c) any device that was also found to be substantially equivalent under the 510(k) pathway (FDA 2018a). Substantial equivalence, in turn, occurs when the device has the same intended use and technological characteristics, or has different technological characteristics but is as safe and effective as the predicate. The FDA determines whether the device is as safe and effective as the predicate device by reviewing the scientific methods used to evaluate differences in technological characteristics and performance data (FDA 2019).
@@
The approach used by the FDA has been heavily criticized. For example, the National Academy of Medicine stated that “the 510(k) process cannot be transformed into a premarket evaluation of safety and effectiveness as long as the standard for clearance is substantial equivalence to any previously cleared device” (FDA 2022b). In 1996, the U.S. Supreme Court also concluded that “the 510(k) process is focused on equivalence, not safety.” (Lohr vs Medtronic 1996). More recently, the FDA has also recognized the shortcomings of its 510(k) pathway indicating the need to add alternative options to demonstrate that a device is as safe and effective as a legally marketed device (Challoner and Senate 2011).
@@
Several studies have also highlighted the potential drawbacks associated with the FDA’s 510(k) pathway, indicating that 71% of devices recalled by the FDA for safety concerns between 2005 and 2009 had initially been cleared through the 510(k) pathway (Zuckerman et al. 2011), or that 11% of devices cleared via the 510(k) pathway between 2003 and 2018 had later been subject to Class I or II recalls (Everhart et al. 2023). Despite all these concerns, there is currently a clear lack of evidenced-based understanding on how the FDA can improve its 510(k) pathway. This is evidenced by the FDA’s 2023 solicitation for feedback (FDA 2023b and FDA 2023f) and its 2018 Medical Device Safety Action Plan for modernizing the 510(k) pathway (FDA 2018b). Our main goal in this paper is to develop a data-driven approach that can assist FDA in improving its 501(k) pathway. To this end, we note that developing a highly accurate predictive model of recall risk can enable the FDA to pinpoint devices that might require more rigorous examination or necessitate stronger evidential support before approval, ultimately helping the FDA to mitigate unnecessary harm to patients. Developing such a prediction model requires access to large data containing comprehensive information on both applicant devices and predicates as well as detailed data engineering processes to generate useful insights. This level of detail cannot be directly obtained using publicly available FDA datasets and requires significant additional data collection and pre-processing efforts.
@@
We also note that just developing an accurate prediction model using large-scale data might not be sufficient to develop a policy for improving the FDA’s 510(k) pathway. The first challenge is related to the nuanced differences of some 510(k) devices, which may make it hard for a policy fully relying on a machine learning (ML) model to be effective. There is a vast literature in support of the interaction between ML models and humans to achieve better overall performance. Thus, we hypothesize that a combined human-algorithm approach to evaluating medical devices can significantly improve FDA’s 510(k) pathway. The second challenge is that any such policy must also consider the limited recourses available to FDA to perform further examination when a device is deemed risky for approval by the model. The FDA receives a high volume of 510(k) submissions each year and has a goal of making a decision within 90 days. The sheer number of submissions can strain the FDA’s resources, resulting in a potential backlog and delays in the review process. This backlog may limit the depth and rigor of the review, potentially impacting the thoroughness of the evaluation and the ability to identify potential risks or issues for a variety of devices. Thus, to rigorously improve the 510(k) pathway, one needs to also develop an approval and evaluation policy guideline that not only can benefit from the predicted risk of recall, but can also take into account FDA’s limited resources for further evaluation of devices that are not a clear-cut for approval or rejection decisions based on the model’s predicted risk.
@@ 1.1 Overview and Contributions
We introduce a data-driven clearance policy aimed at assisting the FDA in improving its 510(k) pathway. Our proposed policy creates a combined human-algorithm approach by deferring some decisions to human exerts and others to a well-trained ML model. Specifically, our policy generates recommendations for direct approval, rejection, or further evaluation by human experts using a variety of the applicant device’s characteristics as well as those of its predicates. Our work is based on close corroboration and extensive discussions with a collaborator (and co-author of this paper) who has been directly involved in numerous related FDA regulatory improvement projects.
@@
Our approach consists of two main steps. In the first step, we collect and assemble FDA data related to 510(k) devices, and develop ML models capable of estimating each device’s risk of recall based on the information available at the time of submission. Specifically, employing a text-scraping algorithm, we start by extracting information regarding both applicant devices and their corresponding predicates, including their characteristics and recall information. Through extensive discussions with our collaborator, we also conduct thorough data engineering to create variables that aid in predicting a recall event. Building on our analysis, we then train and test several ML models and pick the best one, which achieves a cross-validation Area Under the Curve (AUC) score of 0.77. Our findings using this ML model suggest that the number of recalls reported for predicate devices, along with variables relevant to the age of the predicates, hold significant predictive power for a recall event. This validates speculations and initial empirical findings in the recent literature but through an ML lens that is trained on large-scale data involving over 31,000 medical devices produced by 12,000 manufactures from over 65 countries. Interestingly, however, unlike these more known facts, our analysis indicates that the timing of recalls reported for predicates of an applicant device and the age of the latest-approved predicate provide valuable information for predicting a recall event.
@@
In the second step, utilizing the ML model for risk assessment, we proceed to develop a data-driven clearance policy that can assist the FDA in decision-making. To this end, we make use of an optimization approach that sets decision-making thresholds. Our approach takes into account the existence of ranges within the risk spectrum where the predicted risk is not informative enough to make a recommendation, necessitating human expert attention. These thresholds are strategically determined to balance the precision of decision-making with the workload burden imposed on the FDA during the decision-making process. The primary computational challenge lies in enforcing a workload constraint for the FDA, turning the optimization model into a non-convex optimization problem. To address this, we develop an algorithm that provides an approximation by making use of Lagrangian relaxation. This algorithm efficiently identifies effective thresholds on predicted recall risks through an iterative solution procedure. We also emphasize that our policy is designed to be capable of providing additional information (risk labels) for applicant devices that require further evaluation by human experts (e.g., by an FDA committee).
@@
We investigated the potential benefits of our policy compared to the current practice of the FDA using a unique large-scaled dataset that we assembled based on the public FDA datasets (Section 5). Our results indicated a 38.9% improvement in the recall rate and a 43.0% reduction in the FDA’s workload. Additionally, we estimated the potential cost-savings from implementing our data-driven policy through estimating medical device replacement costs by medical specialties (Section 5.1). For the very first time, this is done by carefully determining the medical specialty for over 99% of the 1,351 unique Healthcare Common Procedure Coding System (HCPCS) codes/descriptions for the CMS administrative claims data during 2013-2020. Our analysis indicate significant potential annual cost-savings from implementing our policy ranging between \$2.4 billion to \$2.7 billion, as well as a notable decrease in adverse event outcomes and an enhancement in patient safety.
@@
Our results provide valuable insights for the FDA and address main concerns about the current 510(k) process (Section 6). First, we observe that while there are some easy cases suitable for direct algorithmic decision-making, some other cases are difficult and require more in-depth evaluation and human judgment. This observation confirms our hypothesis that there is a need for a combined human-algorithm approach to integrate the FDA’s expertise with quantitative evidence. Next, our results address the FDA’s concerns communicated in its recent solicitation announcements as well as its goal to establish best practices to evaluate the safety of medical devices (FDA 2023b). Specifically, we find that best practices should take advantage of the fact that the number of recall events for predicates, particularly recent recalls, is a significant predictor of recall risk for an applicant device. Furthermore, we observe that the age of the latest-approved and earliest-approved predicates, as well as the average age of the predicates are among other important predictors of recall risk. Finally, our results contribute to the recent call for feedback by the FDA on the opportunities and challenges of using artificial intelligence (AI) and ML in the development of drugs and medical devices (FDA 2023f). Our work highlights the importance of crucial considerations in the context of utilizing AI/ML to enhance the evaluation of medical devices, and highlights the importance of employing a combined human-algorithm approach.
@@ 1.2 Literature Review
Medical Device Recalls. There is a vast literature focusing on comparing the risk of recall for devices that received approval via the 510(k) pathway and those that received approval via premarket approval (see, e.g., Day et al. 2016, Connor et al. 2017, Janetos et al. 2017, Talati et al. 2018, and Dubin et al. 2021). There are also a few empirical studies (see, e.g., Wowak et al. 2021, Ball et al. 2018, Wowak et al. 2021, and Ball et al. 2018) related to our work that examine factors influencing recall-related events of medical devices. Another related set of studies investigate the relationship between the characteristics of predicate medical devices and the recall events of 510(k) devices (Everhart et al. 2023 and Kadakia et al. 2023). The study of Mukherjee and Sinha 2018, presents a predictive model for risk of recall using the adverse events occurring after the approval time. However, predictive models focusing on estimating the recall risk of an applicant device based on the information available at the time of submission remain scarce. In our study, we address this gap by developing a recall risk prediction model and recommending a clearance policy for the FDA, wherein the policy’s recommendations are constrained by the available information at the time of the 510(k) application submission as well as the workload that can be imposed to FDA for further evaluations by human experts prior to decision-making.
@@
Medical Diagnostic Decision. Various studies in the literature focus on medical diagnostic decisions with the goal of optimizing pre-diagnosis or follow-up choices (see, e.g., Ayvaci et al. 2012, Zhang et al. 2012, and Bayati et al. 2018). Our work is closest to the subset of the literature that aims at incorporating a decision support tool in the diagnostic decision and potential biases in the process (see, e.g., Ahsen et al. 2019 and Jussupow et al. 2021). Our work is also closely related to studies that aim at improving quality versus speed trade-offs in diagnostic and triage decisions (see, e.g., Saghafian et al. 2018). Our work differs form this stream of literature as it develops a data-driven policy for improving the FDA’s 510(k) pathway in which both ML based predictions and workload considerations are taken into account.
@@
Threshold-Based Decision-Making. There are several existing methods to utilize risk predictions for decision-making by determining a single decision threshold, ranging from utility-based methods (see e.g., Jund et al. 2005, Felder and Mayrhofer 2014, and van Giessen et al. 2018) to receiver operating characteristic (ROC) method (see, e.g., Hajian-Tilaki 2013 and Hong et al. 2021) and Bayesian decision theory (see, e.g., Sheppard and Kaufman 2005 and Weise et al. 2006). Our works is closest to the subset of literature focusing on three-way classifications with three categories of positive, negative, and undecided based on the evidence (see, e.g., Si et al. 2017, Yao 2010, Yao and Zhou 2016, and Garcia et al. 2020). However, our work differs from these studies in two major ways. First, balancing workload plays a significant role in our work. In a three-way classification using two decision thresholds, setting the threshold conservatively often results in achieving a high performance measure in the positive and negative regions. However, it may lead to assigning many other instances to the undecided region which often will require attention from human experts. Consequently, this approach may not significantly reduce the workload in the system, though balancing the workload (speed) and quality of the decisions made is a crucial element in many systems (see, e.g., Saghafian et al. 2018). Second, the existing methods for finding the decision thresholds without workload considerations are less complex and often can be solved using simple heuristics (Si et al. 2017), Bayesian rough sets (Yao 2010, Yao and Zhou 2016), or a linear program (Garcia et al. 2020). In contrast, our methodology requires different techniques to handle the non-linearity in optimizing decision thresholds. Our solution technique, obtained by deriving structural properties of our optimization model and introducing a Lagrangian-based algorithm, allows us to capture the complexity introduced by workload considerations and identify the decision thresholds efficiently.
@@
Human-in-the-Loop Approaches. An expanding body of research suggests that ML models can outperform humans in making predictions across a wide range of domains (see, e.g., Liu et al. 2018, Shen et al. 2019, Boloori et al. 2022, Ang et al. 2022). While the state-of-the-art ML models exhibit impressive performance, in high-risk domains such as healthcare, there is reluctance to fully embrace automated AI systems and eliminate humans entirely from the loop due to the inherent distrust in AI systems and lack of robustness (Association et al. 2019). Recent literature has highlighted the benefits of incorporating human judgment into the ML model deployment process, leading to the development of human-in-the-loop approaches. The idea of keeping humans in the loop has been implemented in various ways. There have been attempts to design interactive and active ML systems that continuously learn from humans (Wu et al. 2022) or even go beyond human-in-loop mechanisms by systematically incorporating symbiotic learning (Muller 2022, Saghafian 2023). Other attempts include combining separate human and algorithm outputs (e.g., Blattberg and Hoch 1990, Goodwin 2000), introducing systems to elicit human judgment for prediction algorithms (Ibrahim et al. 2021), and learning the human experts’ intuition for risk prediction (Orfanoudaki et al. 2022). Another approach to keeping humans in the loop is AI augmentation, where the main idea is to have AI systems work alongside humans and collaborate with them. This idea is different from automation that results in replacing humans with AI (Daugherty and Wilson 2018, Miller 2018). Our work provides further evidence on the latter approach, providing a regulatory policy that allows human experts to concentrate on complex cases with the assistance of AI/ML.
@@
The reminder of the paper is organized as follows. Section 2 provides a summary of our data collection and pre-processing steps. Section 3 describes the development and evaluation of prediction models. Section 4 introduces our data-driven clearance policy, its analytical properties, and our solution methodology. Section 5 provides our empirical results. Finally, we present managerial insights and concluding remarks in Sections 6 and 7, respectively.

### v2

@@ 1 Introduction
The medical device approval pathway is a function of the level of control necessary to provide reasonable assurance of a device’s safety and effectiveness (FDA 2018a). In general, devices posing a greater degree of risk are denoted by class and face greater regulatory controls. Although there are four common types of approval pathways (FDA 2022a), the vast majority of medical devices are cleared under the Premarket Notification 510(k) pathway. For the year 2022, over 3,000 devices were cleared under the 510(k) pathway (FDA 2023a), while less than two hundred devices were cleared under the Humanitarian Device Exemption, Premarket Approval, and De Novo pathways combined (FDA 2023e, FDA 2023d, and FDA 2023c).
@@
The 510(k) pathway was developed both to reduce the burden for device manufacturers bringing medium-to-low risk (Class II and I) devices to market and to address the limited resources of the United States Food and Drug Administration (FDA) (Kramer and Yeh 2023). Devices cleared under the 510(k) pathway must demonstrate that the device is “substantially equivalent” to a legally marketed device, commonly referred to as a predicate device (FDA 2022b). A predicate device can be (a) any device that is legally marketed prior to May 28, 1976 (a preamendments device), for which clinical testing was not required, (b) any device that was reclassified from Class III (high risk) to Class II or Class I (medium-to-low-risk), or (c) any device that was also found to be substantially equivalent under the 510(k) pathway (FDA 2018a). An applicant seeking FDA clearance for a device through the 510(k) pathway is required to select at least one predicate but may choose multiple predicates at their discretion. The FDA provides guidance on how to find and effectively use predicate devices. In general, manufacturers are asked to identify a predicate that is most similar to the device under review with respect to indications for use and technological characteristics (FDA 2018b). However, manufacturers may choose to reference multiple predicates when combining features from different devices with the same intended use, seeking clearance for a device with multiple intended uses, or requesting multiple indications for use under a shared intended use (FDA 2014). The applicant is not required to identify the universe of predicate devices that meet this standard. Substantial equivalence, in turn, occurs when the device has the same intended use and technological characteristics, or has different technological characteristics but is as safe and effective as the predicate. The FDA determines whether the device is as safe and effective as the predicate device by reviewing the scientific methods used to evaluate differences in technological characteristics and performance data (FDA 2019).
@@
The 510(k) pathway remains the principal route through which most moderate-to-low risk devices reach the U.S. market and offers the advantage of providing patients with timely access to incremental innovation. Nonetheless, the pathway has also attracted considerable scrutiny. For example, the National Academy of Medicine stated that “the 510(k) process cannot be transformed into a premarket evaluation of safety and effectiveness as long as the standard for clearance is substantial equivalence to any previously cleared device” (FDA 2022b). In 1996, the U.S. Supreme Court also concluded that “the 510(k) process is focused on equivalence, not safety.” (Lohr vs Medtronic 1996). More recently, the FDA has also recognized the shortcomings of its 510(k) pathway indicating the need to add alternative options to demonstrate that a device is as safe and effective as a legally marketed device (Challoner and Senate 2011).
@@
The FDA classifies an action as a recall when the manufacturer corrects or removes a device that is out of compliance with FDA regulations. Typically, a recall event occurs because the device is faulty, poses a health hazard, or both. A recall may involve a correction that allows the device to remain in use, or a removal that requires physical return or destruction of affected units (FDA 2018e). The 510(k) clearance is not automatically revoked when a recall occurs; instead, the FDA may require additional data, restrict distribution, or, in some cases, withdrawing the device from the market altogether. A single device can accumulate multiple recall events over its commercial life. Several studies have also highlighted the potential drawbacks associated with the FDA’s 510(k) pathway, indicating that 71% of devices recalled by the FDA for safety concerns between 2005 and 2009 had initially been cleared through the 510(k) pathway (Zuckerman et al. 2011), or that 11% of devices cleared via the 510(k) pathway between 2003 and 2018 had later been subject to Class I or II recalls (Everhart et al. 2023). Despite all these concerns, there is currently a clear lack of evidence-based understanding on how the FDA can improve its 510(k) pathway. This is evidenced by the FDA’s 2023 solicitation for feedback (FDA 2023b and FDA 2023f) and its 2018 Medical Device Safety Action Plan for modernizing the 510(k) pathway (FDA 2018c). Our main goal in this paper is to develop a data-driven approach that can assist the FDA in improving its 510(k) pathway. To this end, we note that developing a highly accurate predictive model of recall risk can enable the FDA to pinpoint devices that might require more rigorous examination or necessitate stronger evidential support before approval, ultimately helping the FDA to mitigate unnecessary harm to patients. Developing such a prediction model requires access to large data containing comprehensive information on both applicant devices and predicates as well as detailed data engineering processes to generate useful insights. This level of detail cannot be directly obtained using publicly available FDA datasets and requires significant additional data collection and pre-processing efforts.
@@
We also note that just developing an accurate prediction model using large-scale data might not be sufficient to develop a policy for improving the FDA’s 510(k) pathway. The first challenge is related to the nuanced differences of some 510(k) devices, which may make it hard for a policy fully relying on a machine learning (ML) model to be effective. There is a vast literature in support of the interaction between ML models and humans to achieve better overall performance. Thus, we hypothesize that a combined human-algorithm approach to evaluating medical devices can significantly improve FDA’s 510(k) pathway. The second challenge is that any such policy must also consider the limited resources available to the FDA to perform further examination when a device is deemed risky for approval by the model. The FDA receives a high volume of 510(k) submissions each year and has the goal of making a decision within 90 days. The sheer number of submissions can strain the FDA’s resources, resulting in a potential backlog and delays in the review process. This backlog may limit the depth and rigor of the review, potentially impacting the thoroughness of the evaluation and the ability to identify potential risks or issues for a variety of devices. Thus, to rigorously improve the 510(k) pathway, one needs to also develop an approval and evaluation policy guideline that not only can benefit from the predicted risk of recall, but can also take into account FDA’s limited resources for further evaluation of devices that are not a clear-cut for approval or rejection decisions based on the model’s predicted risk.
@@ 1.1 Overview and Contributions
We introduce a data-driven clearance policy aimed at assisting the FDA in improving its 510(k) pathway. Our proposed policy creates a combined human-algorithm approach by deferring some decisions to human experts and others to a well-trained ML model. Specifically, our policy generates recommendations for direct approval, rejection, or further evaluation by human experts using a variety of the applicant device’s characteristics as well as those of its predicates. Our work is based on close corroboration and extensive discussions with a collaborator (and co-author of this paper) who has been directly involved in numerous related FDA regulatory improvement projects.
@@
Our approach consists of two main steps. In the first step, we collect and assemble FDA data related to 510(k) devices, and develop ML models capable of estimating each device’s risk of recall based on the information available at the time of submission. Specifically, employing a text mining algorithm, we start by extracting information regarding both applicant devices and their corresponding predicates, including their characteristics and recall information. Through extensive discussions with our collaborator, we also conduct thorough data engineering to create variables that aid in predicting a recall event. Building on our analysis, we then train and test several ML models and pick the best one, which achieves a cross-validation Area Under the Curve (AUC) score of 0.78. Our findings using this ML model suggest that the number of recalls reported for predicate devices, along with variables relevant to the age of the predicates, hold significant predictive power for a recall event. This validates speculations and initial empirical findings in the recent literature but through an ML lens that is trained on large-scale data involving over 31,000 medical devices produced by 12,000 manufacturers from over 65 countries. Interestingly, however, unlike these more known facts, our analysis indicates that the timing of recalls reported for predicates of an applicant device and the age of the predicates provide valuable information for predicting a recall event.
@@
In the second step, utilizing the ML model for risk assessment, we proceed to develop a data-driven clearance policy that can assist the FDA in decision-making. To this end, we make use of an optimization approach that sets decision-making thresholds. Our approach takes into account the existence of ranges within the risk spectrum where the predicted risk is not informative enough to make a recommendation, necessitating human expert attention. These thresholds are strategically determined to balance the precision of decision-making with the workload burden imposed on the FDA during the decision-making process. The primary computational challenge lies in enforcing a workload constraint for the FDA, turning the optimization model into a non-convex optimization problem. To address this, we develop a nested search algorithm that provides an approximation by leveraging the structural properties of our problem. This algorithm efficiently identifies effective thresholds on predicted recall risks through an iterative solution procedure. We also emphasize that our policy is designed to be capable of providing additional information for applicant devices that require further evaluation by human experts (e.g., by an FDA committee).
@@
We investigated the potential benefits of our policy compared to the current practice of the FDA using a unique large-scaled dataset that we assembled based on the public FDA datasets (Section 5). Our results indicated a 32.9% improvement in the recall rate and a 40.5% reduction in the FDA’s workload. Additionally, we estimated the potential cost savings from implementing our data-driven policy through estimating medical device replacement costs by medical specialties (Section 5.1). For the very first time, this is done by carefully determining the medical specialty for over 99% of the 1,351 unique Healthcare Common Procedure Coding System (HCPCS) codes/descriptions for the CMS administrative claims data during 2013-2020. Our analysis indicates significant potential annual cost savings of \$1.7 billion from implementing our policy.
@@
Our results provide valuable insights for the FDA and address main concerns about the current 510(k) process (Section 6). First, we observe that while there are some easy cases suitable for direct algorithmic decision-making, some other cases are difficult and require more in-depth evaluation and human judgment. This observation confirms our hypothesis that there is a need for a combined human-algorithm approach to integrate the FDA’s expertise with quantitative evidence. Next, our results address the FDA’s concerns communicated in its recent solicitation announcements as well as its goal to establish best practices to evaluate the safety of medical devices (FDA 2023b). Specifically, we find that best practices should take advantage of the fact that the number of recall events for predicates, particularly recent recalls, is a significant predictor of recall risk for an applicant device. Furthermore, we observe that the age of the predicates is among other important predictors of recall risk. Finally, our results contribute to the recent call for feedback by the FDA on the opportunities and challenges of using artificial intelligence (AI) and ML in the development of drugs and medical devices (FDA 2023f). Our work highlights the importance of crucial considerations in the context of utilizing AI/ML to enhance the evaluation of medical devices, and highlights the importance of employing a combined human-algorithm approach.
@@ 1.2 Literature Review
Medical Device Recalls. There is a vast literature focusing on comparing the risk of recall for devices that received approval via the 510(k) pathway and those that received approval via premarket approval (see, e.g., Day et al. 2016, Connor et al. 2017, Janetos et al. 2017, Talati et al. 2018, and Dubin et al. 2021). There are also a few empirical studies (see, e.g., Wowak et al. 2021 and Ball et al. 2018) related to our work that examine factors influencing recall-related events of medical devices. Another related set of studies investigates the relationship between the characteristics of predicate medical devices and the recall events of 510(k) devices (Everhart et al. 2023 and Kadakia et al. 2023). The study of Mukherjee and Sinha 2018 presents a predictive model for the risk of recall using the adverse events occurring after the approval time. However, predictive models focusing on estimating the recall risk of an applicant device based on the information available at the time of submission remain scarce. In our study, we address this gap by developing a recall risk prediction model and recommending a clearance policy for the FDA, wherein the policy’s recommendations are constrained by the available information at the time of the 510(k) application submission as well as the workload that can be imposed to the FDA for further evaluations by human experts prior to decision-making.
@@
To our knowledge, there are no previous scholarly publications that have systematically assessed the costs of recalling the spectrum of medical devices. Our review of the literature uncovered two notable publications on this topic. The first publication was a 2017 report by the Office of Inspector General (OIG). Consistent with our own literature review, this report noted that there is no reliable up-to-date estimate of Medicare costs associated with recalled medical devices (HHS 2017). The second publication was a white paper published by the McKinsey Center for Government and cited by the FDA in a presentation discussing the benefit of reducing medical device failure cost (Tack 2021). To address this evidence gap, we assembled a dataset based on the administrative claims data titled “Medicare Durable Medical Equipment, Devices & Supplies” (MDMEDS). This dataset allows us to estimate cost savings for payers, which are quantified by the avoidance of replacement costs associated with device recalls. These savings would result from devices that were cleared by the FDA but would have been rejected by our policy, had it been implemented.
@@
Medical Diagnostic Decision. There is an extensive body of literature on medical diagnostic decisions across both the operations research and machine learning communities (see, e.g., Mirbabaie et al. 2021 and Denton 2023 for broader reviews). Various studies in the literature focus on medical diagnostic decisions with the goal of optimizing pre-diagnosis or follow-up choices (see, e.g., Ayvaci et al. 2012, Zhang et al. 2012, and Bayati et al. 2018). Our work is closest to the subset of the literature that aims at incorporating a decision support tool in the diagnostic decision and potential biases in the process (see, e.g., Ahsen et al. 2019 and Jussupow et al. 2021). Our work is also closely related to studies that aim at improving quality versus speed trade-offs in diagnostic and triage decisions (see, e.g., Saghafian et al. 2018). Our work differs from this stream of literature as it develops a data-driven policy for improving the FDA’s 510(k) pathway in which both ML based predictions and workload considerations are taken into account.
@@
Threshold-Based Decision-Making. There are several existing methods to utilize risk predictions for decision-making by determining a single decision threshold, ranging from utility-based methods (see e.g., Jund et al. 2005, Felder and Mayrhofer 2014, and van Giessen et al. 2018) to the receiver operating characteristic (ROC) method (see, e.g., Hajian-Tilaki 2013 and Hong et al. 2021) and Bayesian decision theory (see, e.g., Sheppard and Kaufman 2005 and Weise et al. 2006). Our works is closest to the subset of literature focusing on three-way classifications with three categories of positive, negative, and undecided based on the evidence (see, e.g., Baram 1998, Si et al. 2017, Yao 2010, Yao and Zhou 2016, Mozannar and Sontag 2020, Garcia et al. 2020). However, our work differs from these studies in two major ways. First, balancing workload plays a significant role in our work. In a three-way classification using two decision thresholds, setting the threshold conservatively often results in achieving a high performance measure in the positive and negative regions. However, it may lead to assigning many other instances to the undecided region which often will require attention from human experts. Consequently, this approach may not significantly reduce the workload in the system, though balancing the workload (speed) and quality of the decisions made is a crucial element in many systems (see, e.g., Saghafian et al. 2018). Second, most of the existing methods for finding the decision thresholds without workload considerations are less complex and often can be solved using simple heuristics (Si et al. 2017), Bayesian rough sets (Yao 2010, Yao and Zhou 2016), or a linear program (Garcia et al. 2020). In contrast, our methodology requires different techniques to handle the non-linearity in optimizing decision thresholds. Our solution technique is obtained by deriving structural properties of our optimization model and introducing a nested search approach.
Human-in-the-Loop Approaches. An expanding body of research suggests that ML models can outperform humans in making predictions across a wide range of domains (see, e.g., Liu et al. 2018, Shen et al. 2019, Boloori et al. 2022, Ang et al. 2022). Typically, there are two main motivations for adopting human-in-the-loop approaches. The first motivation is related to stakeholder trust. While the state-of-the-art ML models exhibit impressive performance, in high-risk domains such as healthcare, there is reluctance to fully embrace automated AI systems and eliminate humans entirely from the loop due to the inherent distrust in AI and lack of robustness (AMA 2019).
@@
The second motivation involves complementary strengths and performance gains. Recent literature has highlighted the benefits of incorporating human judgment into the ML model deployment process, leading to performance gain. The idea of keeping humans in the loop has been implemented in various ways. There have been attempts to design interactive and active ML systems that continuously learn from humans (Wu et al. 2022) or even go beyond human-in-loop mechanisms by systematically incorporating symbiotic learning (Muller 2022, Saghafian 2023). Other attempts include combining separate human and algorithm outputs (Blattberg and Hoch 1990, Goodwin 2000), introducing systems to elicit human judgment for prediction algorithms (Ibrahim et al. 2021), and learning the human experts’ intuition for risk prediction (Orfanoudaki et al. 2022). Another approach to keeping humans in the loop is AI augmentation, where the main idea is to have AI systems work alongside humans and collaborate with them. This idea is different from automation that results in replacing humans with AI (Daugherty and Wilson 2018, Miller 2018). Our work provides further evidence on the latter approach, providing a regulatory policy that allows human experts to concentrate on complex cases with the assistance of AI/ML.
@@
The remainder of the paper is organized as follows. Section 2 provides a summary of our data collection and pre-processing steps. Section 3 describes the development and evaluation of prediction models. Section 4 introduces our data-driven clearance policy, its analytical properties, and our solution methodology. Section 5 provides our empirical results. Finally, we present managerial insights and concluding remarks in Section 6.